Hi, I don't get this can you please help me.
The question:
If a x b = 2, b x c = 24 and a x c = 3 ( a, b, c are all positive numbers) what is the value of a + b + c?
I thought that 1 x 2 = , 2 x 12 = 24 but 1 x 12 is not 3 so know I am really stuck
Thanks Olga

Answer by mananth(16946)

If a x b = 2, b x c = 24 and a x c = 3 ( a, b, c are all positive numbers)

a x b = 2, b x c = 24
divide
a/c = 2/24 =1/12
a*c =3
multiply both
a^2=1/4
a=1/2
a*b =2
1/2 *b=2
b=4
a*c=3
1/2 *c=3
c=6
find a+b+c